OpenStudy (z4k4r1y4):

Show that x=Be^(µt) is a solution of mx ̈+kx=0

OpenStudy (irishboy123):

firstly, try it as a solution if: x = B e^(µt) then: x' = µ B e^(µt) and: x'' = µ^2 B e^(µt) so mx'' + kx = B e^(µt) { m µ^2 + k } = 0 does that look like a solution, or is more needed? m µ^2 + k = 0 is clearly unhelpful